OpenStudy (needhelpp101):

.) Express the repeating decimal number as a quotient of two integers. 28. --- 81 A.3125/11 B. 83/11 C.317/11 D.5/52

OpenStudy (anonymous):

Is it 28.81

OpenStudy (needhelpp101):

yes

OpenStudy (anonymous):

x = 28.81* 100x = 2881.81* 99x = 2853 x = 2853/99 = 317/11

OpenStudy (anonymous):

Let x=28.818181.... \[100x=2881.81....\] \[99x=2881.81....-28.818181.....\] \[99x=2853\] \[x=\frac{317}{11}\]

OpenStudy (anonymous):

Dang you got there first lolz. too slow.

OpenStudy (needhelpp101):

Ty Karan and Azteck :D <3